OpenStudy (anonymous):

8-16). Solve the following equation by applying the squaring property of equality. Be sure to check all solutions in the original equation. (Enter solutions from smallest to largest. sqrt(2 x + 1) - 1 = 4

OpenStudy (anonymous):

sqrt(2x+1)-1=4 sqrt(2x+1)=5 2x+1=25 2x=24 x=12